# Two-point correlators in de Sitter-prepared states with bra-ket wormholes

### v1

@@ 4.3 Observable 1 - Comparing longer and shorter inflation
Figure 4: Observable 1: $\langle\phi_{k}\phi_{-k}\rangle$ (left) and $\langle\phi(\theta)\phi(0)\rangle$ (right), compared between different $k_{\rm max}$-universes. Wormhole contributions are visible as deviation from HH results (dashed) for $k_{\rm max}\gtrsim 60=k_{\rm max}^{c}$. Small-$k$ or long-range correlations are enhanced.
@@
Observable 1 is to compare dS prepared states of different universes with different $k_{\rm max}$, including all modes up to $k_{\rm max}$. Questions for Observable 1: Given the gravitational scale $S_{\rm dS}$, is phase transition and its internal time scale consistent with information physics and chaotic nature of horizon? Here, we attempt to interpret $k_{\rm max}$ as some kind of time internal to dS, with aforementioned caution.
Two-point correlations in the full state are calculated by Eq. (42) and (43). We use $S_{\rm dS}=76,\tau_{0}=0.01$ and $k_{\rm max}=150,k_{\rm min}=0,m=0.3H$, if not mentioned otherwise.
@@
Above all, Fig. 4 demonstrates the phase transition mechanism. Wormhole effects on $\langle\phi_{k}\phi_{-k}\rangle={\rm Tr}[\rho\,\phi_{k}\phi_{-k}]$ are visible (as deviations from HH expected results) only for universes with large enough $k_{\rm max}\gtrsim 60$. Once the state $\rho$ is dominated by wormhole effects with large $k_{\rm max}$, the deviation is visible predominantly through small-$k$ enhancement of correlation. These are as expected by the mechanism in Sec. 4.2. The fact that wormholes dominate only after long duration of dS stage may be consistent with its possible role of resolving information problem that appears late.
@@
Fig. 4 also makes a consistency check that large-$k$ mode correlations always agree with the leading-order HH results. It is not that the HH term (first term) in Eq. (42) dominates the wormhole(WH) term (second term) for large-$k$ correlations, but that ${\rm Im}A_{k}^{WH}\simeq{\rm Im}A_{k}^{HH}$ for large $k$ in Eq. (46) so that even for WH-dominated states large-$k$ correlation is always that of HH result. This is actually an interesting property of mode functions that satisfy this consistency relation at large $k$ while allowing wormhole dominant solutions in other regimes.
@@
Fig. 4 right panel shows real-space correlations. Unlike scale invariant HH results, WH-dominated spatial correlation is largest between antipodal points $\theta=\pi$, while rapidly decreasing for short-distance correlation $\theta\to 0,2\pi$ (but still larger than HH results). This enhancement of long-range correlation was expected if WH solution is to be consistent with KSW. And this might be attributed to wormhole’s stereotypical property of entangling widely separated regions of space or shortening a geodesic.
@@
Figure 5: Observable 1: Correlation does not decay forever, but ramps up and reaches a plateau, due to bra-ket wormhole contributions. This feature seems to be consistent with the finiteness of dS horizon entropy. Quasinormal decays of superhorizon modes may reflect the chaotic nature of dS.
@@
Fig. 5 shows an important result: correlation does not decay indefinitely with time $k_{\rm max}$, but ramps up and plateaus at late time. Similar behavior was found in the AdS-black hole system Saad:2018bqo; Saad:2019lba; Saad:2019pqd, which was largely expected based on the finiteness of Hilbert space Maldacena:2001kr; Witten:2001kn; Goheer:2002vf. It was also expected in dS Aalsma:2022eru; Chapman:2022mqd; Mirbabayi:2023vgl.
@@
As a bonus, Fig. 5 also exhibits quasinormal decays of superhorizon modes (from horizon exit until the end of inflation), shown as the slow decrease of HH results with $k_{\rm max}$. Technically, this effect is incorporated by $(2\pi/\ell)^{-\Delta_{-}}=k_{\rm max}^{-\Delta_{-}}$ factor in Eq. (3(b)). This exponential decay is originated from the exponential expansion of dS universe, but it also suggests that dS is a chaotic quantum system Shenker:2013pqa; Maldacena:2015waa; Aalsma:2020aib.
Figure 6: $k_{\rm max}$ at which bra-ket wormhole begins to dominate the total partition function. It turns out to scale as $\propto S_{\rm dS}$ ($Ht\propto\log S_{\rm dS}$) for fixed $\tau_{0}=0.01$, reminiscent of the fast scrambling physics. But exact behavior depends on wormhole stabilization mechanism.
The timing of WH dominance in our solution seems to follow the fast scrambling time $Ht\propto\log S_{\rm dS}$, for fixed $\tau_{0}=0.01$, as shown in Fig. 6. Different fixed value of $\tau_{0}=0.001$ also leads to similar behavior (Fig. 12). But true wormhole stabilization shall determine $\tau_{0}$ in relation to $S_{\rm dS}$, then this behavior may change. The exact dependence is not known. It is generally expected that distinguishing a mixed state from a pure state in the Hilbert space of dimension $e^{S}$ requires to measure $S$ entropies Page:1993wv; Arkani-Hamed:2007ryv. This typical time is the Page time $t\sim\frac{1}{H}\cdot S$. But for a chaotic system, information can leak out from much earlier time, so called the fast scrambling time $Ht\sim\log S_{\rm dS}$ Hayden:2007cs; Maldacena:2015waa. There are even arguments that dS is a hyperfast scrambling system Sekino:2008he; Susskind:2011ap; Susskind:2021esx. Thus, the timing that falls into this range is an interesting result that we do not yet understand why so.
@@ 4.4 Observable 2 - CMB-like correlation after reheating
Observable 2 is CMB-like two-point correlations measured by an observer in the flat (or FRW) space after reheating. At each time of measurement (parameterized by $k_{\rm min}$), we combine only $k_{\rm min}\leq k\leq k_{\rm max}$ modes to calculate real-space correlations in Eq. (44). $k_{\rm min}$ serves as a time after reheating; decreasing $k_{\rm min}$ means time running forward. $k_{\rm max}$ is fixed, and the dS prepared state is fixed. This observable is what a post-reheating observer in the given universe would really measure as a function of his cosmological time, via CMB or large scale structures.
@@
Figure 7: Observable 2: Antipodal CMB correlation $\langle\phi(\pi)\phi(0)\rangle$ as a function of time after reheating, $k_{\rm min}$. Wormhole effects visible only at late times seem to be consistent with the observer’s information viewpoint.
@@
Why is this chronological order of CMB observation consistent with information content of the observer? Shortly after reheating, the observer just sees a small amount of information that matches with the HH vacuum, never able to disentangle exponentially suppressed higher topology effects. As he does not gain much information about the fixed state and much uncertainty or possibility remains, the entropy grows. Only after wormhole effects are measured at late times, can he begin to reduce those uncertainties or possibilities, and the entropy decreases. This is also consistent with aforementioned discussion of pure versus mixed state distinction. This chronological phenomenon is very analogous to that of black hole Hawking radiation and the recovery of information(unitarity) after about the Page time.
In a slow-roll inflation, the dS Page time is when the dS horizon entropy is saturated by superhorizon modes generated during inflation. Longer inflation essentially becomes an eternal inflation. What happens to the CMB observation? It was argued that density fluctuations generated during the eternal regime are ${\cal O}(1)$ so that no CMB observation after dS Page time can be made, not violating complementarity Bousso:2006ge; Arkani-Hamed:2007ryv. Our phase transition that occurs much earlier than the Page time (in Observable 1) may cure or change related physics and observations well before the saturation.
@@
How phase transition occurs while being chronologically consistent is notable, slightly different from that in Observable 1. At early times, only a small number of largest-$k$ modes are observed. They by themselves resemble HH, which was a consistency property. It is not that these largest $k$-modes are generated during the HH dominance, rather they are the ones generated at the last stage of dS where wormholes dominate most. But they resemble HH by themselves, even though the fixed state, a sum of the whole modes, is dominated by wormhole effects. As time goes on, smaller $k$-modes re-enter, and discernible wormhole contributions gradually appear Page:1993wv.
@@
(a) Example realization of the CMB field $\phi(\theta)$, without(left) and with(right) wormhole contributions. $k_{\rm max}=150,k_{\rm min}=0,m=0.3H,\tau_{0}=0.01,S_{\rm dS}=76$. Enhanced long-range correlations are clear.
@@
(b) Same as Fig. 9(a), but for heavier $m=0.49H$. Enhancement of long-range correlation is less prominent.
@@
Lastly, Fig. 9(a) and 9(b) show example late-time CMB distributions on the 1-dimensional space; full results on the right panel is compared with results without wormholes on the left panel. In these sample realizations, field values $\phi_{k}$ are random selected according to the Gaussian probability with the width given by $\langle\phi_{k}\phi_{-k}\rangle$. The wormhole effects show up as stronger long-range correlations as expected. This can also be measured as enhanced dipole anisotropy. It is more prominent for light scalars in Fig. 9(a).
@@ 5 Summary and Discussion
In Sec. 4, we have introduced two correlation observables of the dS prepared state by utilizing two time variables. The relevant parameter spaces they are respectively probing and utilizing are summarized in the full parameter space $\{k_{\rm max},k_{\rm min}\}$ in Fig. 9. The first observable (along $B$) compares dS prepared states of different universes, allowing to diagnose the dynamics of phase transition into wormhole dominance. The second observable (along $A$ or $A^{\prime}$ depending on which $k_{\rm max}$-universe) is basically CMB observable as a function of observer’s cosmological time after reheating. Each correlation observable was qualitatively consistent by itself, finding decay-ramp-plateau–type behavior, phase transition at about scrambling time, and consistent chronological order. These were possible due to (1) inflationary exit and re-enter, (2) small-$k$ enhancement of correlations, and (3) phase transition mechanism via competition between mode number and topological suppression. Although results are qualitatively consistent, to really understand them in terms of the entropy context and the chaotic nature of scrambling requires to address following limitations.
Figure 10: Contours of $k_{\rm max}-k_{\rm min}^{c}$, the CMB timing that first sees 10% deviation from HH results. $\tau_{0}=0.01$ determines $k^{*}=1/\tau_{0}$ (dotted horizontal) above which all modes look like $H$ states. The wormhole dominance is found in the region above the red boundary $k_{\rm max}^{c}(S_{\rm dS})$. Region AA: Too short inflation does not need wormhole dominance. $B$: Long enough inflation with wormhole-dominant state. $C$, $D$, $E$: limitations of our work; see text for detail.
Stabilization of $\tau_{0}$ with $S_{\rm dS}$.
An important limitation of this work is that $\tau_{0}$ was chosen by hand without relevance to the gravity action (or $S_{\rm dS}$); even with extra fermions, the saddle point is achieved by the balance between fermions and scalars not involving $S_{\rm dS}$. We first discuss problems with fixed $\tau_{0}$, and possible resolutions, and briefly answer how our qualitative conclusions change with $\tau_{0}$ variation.
Fig. 10 shows the contours of $k_{\rm max}-k_{\rm min}^{c}$, the difference of $k_{\rm max}$ and $k_{\rm min}^{c}$, which is the CMB timing of revealing wormhole effects in the $k_{\rm max}$-universe. Along the arrow $A\to B$, universes have longer inflation and larger wormhole effects in the partition function. For universes in $A$, inflation is too short to have/need wormhole dominance. For universes in $B$, our consistent story of wormhole dominance and CMB observation follows. But for universes in $D$, where $k_{\rm max}^{c}<k_{\rm max}<k^{*}=1/\tau_{0}$, early CMB observation would immediately reveal wormhole effects. One may naively expect that since $k_{\rm max}^{c}$ is determined by interplay of $\tau_{0}$ and $S_{\rm dS}$, if $\tau_{0}$ is somehow properly determined in relation to $S_{\rm dS}$ (so not a fixed quantity) the problem of region $D$ might be resolved. But actually in our phase transition mechanism, $k_{\rm max}^{c}$ cannot be larger than $k^{*}=1/\tau_{0}$. Region $D$ cannot be avoided in our mechanism.
But the size of region $D$ may depend on a true wormhole stabilization mechanism. Another missing ingredient might be that due to slow-rolling of the inflaton, $S_{\rm dS}$ becomes maximal toward the end of inflation, making wormhole effects relatively smaller there. And these last modes are the ones that will be observed first by CMB observer. It is also an interesting question by itself how to implement slow-rolling in gravitational path integral, or more generally the response of the gravity action to the matter sector which is absent in the JT model. Lastly, one can make an anthropic argument. No advanced observers can exist right after reheating; it always takes some time for universes to evolve to accommodate advanced observers. Notably, this anthropic argument is unique to dS-flat setup, as gravitating and flat regions are timelike separated; black hole observers can exist simultaneously with the black hole.
In universes with very long inflation (region $C$), the CMB reveal of wormhole is also delayed indefinitely. But there shall be some upper limit on this timing, related to the Page time, so this indefinite delay cannot be correct. This region may be related to eternal inflation. But we do not see clear ways how the fact that inflation is eternal causes problems in this work, and we leave investigation for future works.
In universes in $E$, right to the red boundary, $S_{\rm dS}$ is too large to allow phase transition for any large $k_{\rm max}$. This might also be due to fixed $\tau_{0}$ independently of $S_{\rm dS}$ as well as missing higher-order wormhole saddles and off-shell wormholes Stanford:2020wkf; Jung:2025RMPS. For example, if $\tau_{0}$ saddle were smaller for larger $S_{\rm dS}$, the red boundary would have moved to the right bottom.
We wrap up this issue by asking: under variation of $\tau_{0}$, how would our results change? In Appendix C, we briefly show that for certain choices of parameters $S_{\rm dS},\tau_{0},k_{\rm max}$} related by some scaling, the shape of plots remain similar and our qualitative conclusions still hold. This is especially true for results with fixed $S_{\rm dS}=76$. But for the timing result as a function of $S_{\rm dS}$ do depend on exact stabilization mechanism.

### v2

@@ 4.4 Observable 1 - Comparing longer and shorter inflation
Figure 6: Observable 1: Basic features of wormhole-dominated correlations. $\langle\phi_{k}\phi_{-k}\rangle$ (left) and $\langle\phi(\theta)\phi(0)\rangle$ (right), compared between different $k_{\rm max}$-universes. Wormhole contributions are visible as deviation from HH results (dashed) for $k_{\rm max}\geq k_{\rm max}^{c}\simeq 60$. Small-$k$ or long-range correlations are enhanced with maximum between antipodal points. (left-bottom): the range of each mode-function ratio is set by $\tau_{0}$ for given $S_{\rm dS}$.
@@
Observable 1 is to compare dS prepared states of different universes with different $k_{\rm max}$, including all modes up to $k_{\rm max}$. Two-point correlations in the full state are calculated by Eq. (42) and (43). Questions for Observable 1: Given the gravitational scale $S_{\rm dS}$, is phase transition and its timescale consistent with information physics and the chaotic nature of a horizon? Here, we attempt to interpret $k_{\rm max}$ as some kind of time internal to dS, with aforementioned caution.
@@
Above all, Fig. 6 demonstrates the phase transition mechanism. Wormhole effects on $\langle\phi_{k}\phi_{-k}\rangle={\rm Tr}[\rho\,\phi_{k}\phi_{-k}]$ are visible (as deviations from HH expected results) only for universes with large enough $k_{\rm max}\geq k_{\rm max}^{c}\sim 60$. Once the state $\rho$ is dominated by wormhole effects with large $k_{\rm max}$, the deviation is visible predominantly through enhancements of small-$k$ correlations. These are as expected by the mechanism in Sec. 4.2. The fact that wormholes dominate only after long duration of dS stage may be consistent with its possible role of resolving information problem that appears only late.
@@
Fig. 6 also makes a consistency check that correlations of large-$k$ modes always agree with leading-order HH results. It is not that the HH term (first term) in Eq. (42) dominates the wormhole(WH) term (second term) for large-$k$ correlations, but that ${\rm Im}A_{k}^{WH}\simeq{\rm Im}A_{k}^{HH}$ for large $k>k^{*}\simeq 100(\sim 1/\tau_{0})$ in Eq. (46), so that even for WH-dominated states large-$k$ correlation is always that of HH result.
In the left-bottom panel, the log of each mode-function ratio at $k_{\rm max}=150$ is plotted, showing that the area under the curve (essentially the $Z_{\rm rat}$) scales with $\log Z_{\rm rat}\propto\frac{1}{\tau_{0}}\log\frac{1}{\tau_{0}}$. As $1/\tau_{0}$ sets the scale of $k_{\rm max}^{c}$ and $k_{\rm max}^{*}$, this proves the scaling relation in Eq. (48) and the appearance of the scrambling timescale in Fig. 4.
@@
Fig. 6 right panel shows real-space correlations. Unlike scale invariant HH results, WH-dominated spatial correlation is largest between antipodal points $\theta=\pi$, while rapidly decreasing for short-distance correlation $\theta\to 0,2\pi$ (but still larger than HH results). This enhancement of long-range correlation was expected if WH solution is to be consistent with KSW. And we also argued that this might be attributed to wormhole’s stereotypical property of entangling across the horizon, shortening a geodesic.
@@
Figure 7: Observable 1: Correlation does not decay forever, but ramps up and reaches a plateau, due to bra-ket wormhole contributions. This feature may be consistent with the finiteness of dS horizon entropy. Quasinormal decays of superhorizon modes apparent in the heavy mass case (right panel) may reflect the chaotic nature of dS. Excluding zero modes ($k=0$) results in the overall decrease of correlation, but main features remain.
@@
Fig. 7 shows an important result: correlation does not decay indefinitely with time $k_{\rm max}$, but ramps up and plateaus at late time. Similar behavior was found in the AdS-black hole system Saad:2018bqo; Saad:2019lba; Saad:2019pqd, which was largely expected based on the finiteness of Hilbert space Maldacena:2001kr; Witten:2001kn; Goheer:2002vf. It was expected also in dS Aalsma:2022eru; Chapman:2022mqd; Mirbabayi:2023vgl. Thus, we conclude that this decay-ramp-plateau behavior is an important indication that bra-ket wormholes may really be relevant to the finiteness of dS horizon entropy.
@@
As a bonus, Fig. 7 also exhibits quasinormal decays of superhorizon modes, shown as the slow decrease of HH result, more apparently in the heavy mass case ($m=0.49$). Technically, this effect is incorporated by $(2\pi/\ell)^{-\Delta_{-}}=k_{\rm max}^{-\Delta_{-}}$ factor in Eq. (3(b)). This exponential decay is originated from the exponential expansion of dS universe, but it also suggests that dS is a chaotic quantum system Shenker:2013pqa; Maldacena:2015waa; Aalsma:2020aib.
One concern is about the zero mode ($k=0$) contribution. Formally, there is IR divergence in $m=0$, and large enhancement for finite $m$. Excluding zero modes in Fig. 7 significantly reduces the overall correlation, but main features of phase transition, its timing, and decay-ramp-plateau all remain. Zero modes affect only real-space correlations of Observable 1. Without clear resolution of how to treat them, we are content with keeping zero modes everywhere else.
@@ 4.5 Observable 2 - CMB-like correlation after reheating
Observable 2 is CMB-like two-point correlations measured by an observer in the flat (or FRW) space after reheating. At each time of measurement (parameterized by $k_{\rm min}$), we combine only $k_{\rm min}\leq k\leq k_{\rm max}$ modes to calculate real-space correlations in Eq. (44). $k_{\rm min}$ serves as a time after reheating; decreasing $k_{\rm min}$ means time running forward. $k_{\rm max}$ is fixed, and the dS prepared state is fixed. This observable is what a post-reheating observer in the given universe would really measure as a function of his/her cosmological time, via CMB or large scale structures.
@@
Figure 8: Observable 2: HH+WH as functions of time after reheating $k_{\rm min}$ (decreasing from $k_{\rm max}$) are compared with $H$ alone. Overlaid are irreducible $1\sigma$ error bars from cosmic variance of subhorizon modes at the marked time $k_{\rm min}$. Upper two panels with $(k_{\rm max},k_{\rm min})=(150,80),\,(150,30)$, and lower two with (80,40),(80,10)(80,40),\,(80,10); the upper and lower panels are distinguished by $k_{\rm max}$ versus $k^{*}\simeq 100$. In both cases, early CMB observers would not be able to distinguish wormhole effects.
@@
How phase transition occurs while being chronologically consistent is notable, different from that in Observable 1. Consider first the case with $k_{\rm max}>k^{*}(\simeq 100)$ (the upper panels of Fig. 8). At early times, only a small number of largest-$k$ modes are observed. They by themselves resemble HH, which was a consistency property. It is not that these largest $k$-modes are generated during the HH dominance, rather they are the ones generated at the last stage of dS where wormholes dominate most. But they resemble HH by themselves, even though the fixed state, a sum of the whole modes, is dominated by wormhole effects. As time goes on ($k_{\rm min}$ decreasing), smaller $k$-modes re-enter, and discernible wormhole contributions gradually appear Page:1993wv.
Consider now the case with $k_{\rm max}<k^{*}$ (the lower panels of Fig. 8). In this case, the early CMB modes in principle are different from HH results. But cosmic variance fundamentally limits the early-time distinction of WH+HH from HH results. This is well illustrated in Fig. 8 by $1\sigma$ error bars of cosmic variances. For a simple proxy of the 2D-universe cosmic variance, we take the 4D-universe (2D CMB sky) result for the relative error: $\Delta C_{k}/C_{k}\simeq\sqrt{2/(4k/k_{\rm min}+1)}$, where it was used that the spherical harmonics $\ell$ for the $k$-mode when the $k_{\rm min}$-mode just re-enters the horizon is approximately $\ell(k)\simeq 2k/k_{\rm min}$. Thus, again CMB observers cannot distinguish wormhole effects until late. It is interesting that one form of fundamental constraint—cosmic variance—makes observations consistent with another fundamental physics—CMB complementarity.
Why is this chronological order of CMB observation consistent with information content of the observer? Shortly after reheating, CMB observers just see a small amount of information that match with the HH vacuum, never able to disentangle exponentially suppressed higher topology effects. As the state would essentially look thermal, the entropy grows. Only after wormhole effects are distinguished at late times, the entropy decreases. This order is consistent with the aforementioned discussion of pure versus mixed-state distinction. This chronological phenomenon is also analogous to that of black hole Hawking radiation and the recovery of information(unitarity) only after the Page time.
In a slow-roll inflation, the dS Page time is when the dS horizon entropy is saturated by superhorizon modes generated during inflation. Longer inflation essentially becomes an eternal inflation. What happens to the CMB observation? It was argued that density fluctuations generated during the eternal regime are ${\cal O}(1)$ so that no CMB observations after the dS Page time can be made, not violating complementarity Bousso:2006ge; Arkani-Hamed:2007ryv. Our phase transition that may occur much earlier than the Page time (in Observable 1) may cure or change related physics and observations well before the saturation; but the possibility of Page-time phase transitions are still open in our solution.
A few more minor thoughts on CMB complementarity. One missing ingredient in our calculation might be that $S_{\rm dS}$ in a slow-roll becomes maximal toward the end of inflation, making wormhole effects relatively smaller there. And these last modes are the ones that will be observed at early times by CMB observers. It is also an interesting question by itself how to implement slow-rolling in gravitational path integrals, or more generally the response of the gravity action to the matter sector which is absent in the JT model. One can also make an anthropic argument. No advanced observers can exist right after reheating; it always takes some time for universes to evolve to accommodate advanced observers. Notably, this anthropic argument is unique to dS-flat setup, as gravitating and flat regions are timelike separated; black hole observers can exist simultaneously with the black hole.
@@
(a) Example realization of the CMB field $\phi(\theta)$, without(left) and with(right) wormhole contributions. $k_{\rm max}=150,k_{\rm min}=0,m=0.1,\tau_{0}=0.01,S_{\rm dS}=76$. Enhanced long-range correlations are clear.
@@
(b) Same as Fig. 10(a), but for heavier $m=0.49$. Enhancement of long-range correlation is less prominent.
@@
Lastly, Fig. 10(a) and 10(b) show example late-time CMB distributions on the 1-dimensional space; full results on the right panel is compared with results without wormholes on the left panel. In these sample realizations, field values $\phi_{k}$ are random selected according to the Gaussian probability with the width given by $\langle\phi_{k}\phi_{-k}\rangle$. The wormhole effects show up as stronger long-range correlations as expected. It is more prominent for light scalars. The breaking of scale invariance can also be measured as enhanced dipole anisotropies.
@@ 5 Summary and Discussion
In Sec. 4, we have introduced two correlation observables of the dS prepared state by utilizing two time variables. The parameter spaces they are probing are summarized in the full parameter space $\{k_{\rm max},k_{\rm min}\}$ in Fig. 10. The Observable 1 (along $B$) compares dS prepared states of different universes, allowing to diagnose the dynamics of phase transition into wormhole dominance. The Observable 2 (along $A$ or $A^{\prime}$ depending on which $k_{\rm max}$-universe) is basically the CMB observable as a function of observer’s cosmological time after reheating. Each correlation observable was qualitatively consistent by itself, finding the decay-ramp-plateau–type behavior, phase transitions at about the scrambling time, and consistent chronological orders. These were possible due to (1) inflationary exit and re-enter, (2) small-$k$ enhancement of correlations, (3) phase transition mechanism via competition between mode counting and topological suppression, and (4) irreducible errors from cosmic variance.
A notable result was the possible relevance of scrambling physics to the correlation functions via bra-ket wormholes. We attempted to associate the shock-wave scrambling physics with the shortcut in our bra-ket wormhole geometry. We demonstrated the appearance of the scrambling timescale from the competition between multiplicative mode counting and topological suppression. But Page-time phase transitions are still possible in a small range of the parameter space. The concrete answer and interpretation in terms of entropy and chaotic aspects of dS requires to have a complete wormhole stabilization, which will provide the definite relation of $\tau_{0}$ and $S_{\rm dS}$. If scrambling physics turns out to be correct, our work suggests $1/\tau_{0}\sim S_{\rm dS}$.